OpenStudy (anonymous):

what is cos 45 degrees?

OpenStudy (tkhunny):

Same as the sine of 45º.

OpenStudy (neer2890):

\[\cos(45)=\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (anonymous):